OpenStudy (anonymous):

Help! Pythagorean Theorem in 3-D!!

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

well, what measure do you think would be the base of the triangle?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

it says ts=12 right? so half of that would be the base of the triangle. and since it says that ro=15. so now you have a and b. what is the total of the squared value of them ? (6^2 and 15^2)

OpenStudy (anonymous):

16.2

OpenStudy (anonymous):

yep! thats ypur answer

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

nothing, 16.2 is your answer

OpenStudy (anonymous):

were trying to find "RM"

OpenStudy (anonymous):

i know, you use the pythagorean theorem. (a^2+b^2=c^2) in this case, 15^2 + 6^2 = 16.2^2

OpenStudy (anonymous):

so your answer is 16.2

OpenStudy (anonymous):

Can u help with 1 more?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

first, do 15^2-12^2. then find the square root of that

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok so 9 is the height of the cone. so now that you have the needed factors, plug them into the formula \[\pi r^2\frac{ h }{ 3}\].. so \[\pi*12^2*\frac{ 9 }{ 3}\]

OpenStudy (anonymous):

36.8

OpenStudy (anonymous):

here, to make it easier, do 3.14*144*3. and you dont have to square root it

OpenStudy (anonymous):

1356.4

OpenStudy (anonymous):

youre answer is 1356.5 :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

youre welcome :)

OpenStudy (anonymous):

Are u up for some more questions?

OpenStudy (anonymous):

yes, im doing geometry anyways so im in the mood for it

OpenStudy (anonymous):

find the sqrt of 7^2 and 24^2 and tell me what u get

OpenStudy (anonymous):

25

OpenStudy (anonymous):

thats your answer

OpenStudy (anonymous):

youre going to want to do 19^2-16^2

OpenStudy (anonymous):

C!!

OpenStudy (anonymous):

@tylermcmullen23 u there?

OpenStudy (anonymous):

yep :) c is correct :)

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

no. but close. do 24^2-12^2. you should get 432. sqrt that number

OpenStudy (anonymous):

20.8

OpenStudy (anonymous):

thats your answer :)

OpenStudy (anonymous):

What about this one?

OpenStudy (anonymous):

@tylermcmullen23 u there?

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

try b